OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 6x – 2y = 2 –3x + 4y = 5

OpenStudy (smileyxl3):

12x-4y=4 -3x+4y=5 (-3x+4y=5)+(12x-4y=4)=(9x=9) x=1 Substitute for y

OpenStudy (anonymous):

y=2 Thanks